OpenStudy (anonymous):

how would i estimate the value of √902?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

myininaya (myininaya):

\[\sqrt{900}=\sqrt{9*100}=\sqrt{9}\sqrt{100}=3*10=30\] 31*31=961 so \[\sqrt{902} \] will be closer to 30 then 31. (alot closer)